OpenStudy (chriss):

Compute the product in the given ring: (12)(16) in \[\mathbb{Z}_{24}\] I know the answer is 0 because it's in the book, but what exactly am I being asked here? I tried it as a product of transpositions and didn't find that helpful, then tried just multiplying them as mod 24 but assuming that I'm doing that right, I get 8. So I'm not sure what the problem is actually asking me.

OpenStudy (anonymous):

i dont really know what its asking, but multiplying them mod 24 would give 0 wouldn't it? 12*16 = 12*2*8 = 24*8

OpenStudy (chriss):

yeah ok. I was doing my mod arithmetic incorrectly. Thanks!

OpenStudy (anonymous):

no problem :)